OpenStudy (anonymous):

I'm having trouble cancelling out my units. \[(N \cdot \cancel{m^2}/\cancel{C^2)}\left[\frac{\cancel{C}}{\cancel{m}^2}+\frac{\cancel{C}}{\cancel{m}^2}\right]\] I've N left, but I should have N/C

OpenStudy (jtvatsim):

\[N \cdot \cancel{m^2}/\cancel{C^2)}\left[\frac{\cancel{C}}{\cancel{m}^2}+\frac{\cancel{C}}{\cancel{m}^2}\right]\]

OpenStudy (jtvatsim):

Just reposting the question :)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

nevermind I think I see my mistake, C^2 doesn't cancel out C

OpenStudy (klimenkov):

Very nice. You solved your own problem by yourself.

OpenStudy (klimenkov):

But where is 2 in the numerator?

OpenStudy (anonymous):

2 in the numerator?

OpenStudy (anonymous):

m^2

OpenStudy (klimenkov):

\(\frac{C}{m^2}+\frac{C}{m^2}=\frac{2C}{m^2}\)

OpenStudy (anonymous):

The m's are different, I can't add them. \[8.99\times10^9N\cdot m^2/C^2\left[\frac{(-5\times10^{-6}C)}{(0.11m)^2}+\frac{(5\times10^-6C)}{(0.09m)^2}\right]\]

OpenStudy (anonymous):

I just needed to cancel my units out to make sure I'm doing the calculations correctly, because the answer is supposed to be in N/C

OpenStudy (anonymous):

\[\vec{E}=N/C\]

OpenStudy (klimenkov):

Ah, i understood, that is ok.